OpenStudy (anonymous):

(Composite Functions): if f(x) = x^2 and g(x) = 2x+7, What is f[g(x)] when g(x) = 11

OpenStudy (ranga):

What is f[g(x)] when g(x) = 11 ? Put g(x) = 11 in f[g(x)]. That is, find f(11). f(x) = x^2 f(11) = 11^2 = 121.

OpenStudy (anonymous):

That is incorrect. Solve for g(x) by making x =11. Thus g(x) =2(11) +7 or 29. Now take that solution and plug it into the x value of f(x) = x^2. Thus it would be 29^2 is your answer.